media%2Fba3%2Fba3186cd-7bde-4bd6-b735-a5

Hello,

If you are unable to answers all the questions please can you do only the second set of questions.

Thanks so much!

ORP manufactures one product, cleverly named "Product A". The Income Statement below represents the operating results for the fiscal year just ended, December 31, 2013. NEW JERSEY produced and sold 1,800 tons of Product A during the current year. The manufacturing capacity of NEW JERSEY's facilities is 3,000 tons of Product A. NEW JERSEY CORP INCOME STATEMENT For the year ended December 31, 2017 900,000 Variable costs Manufacturing Selling costs 315,000 180,000 $495,000 405,000 Total variable costs Contribution margin Fixed costs 90,000 112,500 45,000 $247.500 Selling Total fixed costs Net income before income taxes 157,500 Income taxes (40%) Net income after income taxes 94,500 REQUIRED 1. The contribution margin per unit is 2. The contribution margin ratio is 3. The breakeven volume for the year in tons is 4. The breakeven volume in sales revenues is$ 5. By how many tons can sales drop before NEW JERSEY starts to show a loss after taxes? 6. If sales revenue increases by 10% next year, bywhat ercent will NEW JERSEY's pretax income nse? Why is this number greater than 10%?
Answer #1

Hi! As per your suggestion I have solved Ans 5, 6, 7,8.

Please give positive ratings so I can keep answering. Thanks!

Particulars Amount
No. of Units               1,800.00 A
Sales          900,000.00 B
Sales price                  500.00 C=B/A
Variable manufacturing costs          315,000.00 D
Variable manufacturing costs per ton                  175.00 E=D/A
Variable selling costs          180,000.00 F
Variable selling costs per ton                  100.00 G=F/A Bitmap Bitmap Bitmap Bitmap Bitmap Bitmap
Contribution margin          405,000.00 H
Contribution margin per ton                  225.00 I=H/A Bitmap Bitmap Bitmap
Total Variables cost per ton                  275.00 J=E+G
Ans to 5
New Jersey will suffer loss when it will not be able to cover even its fixed costs
Total Fixed costs          247,500.00 K
Contribution margin per ton                  225.00 I
Break even No. of units               1,100.00 L=K/I
Current sales units               1,800.00 A
Difference                  700.00 M=A-L
So Sales should drop by more than 700 tons to get loss after taxes.
Ans to 6
Current Sale price                  500.00 C
New Sale price                  550.00 N=C*110%
Total Variables cost per ton                  275.00 J
New Contribution margin per ton                  275.00 O=N-J
No. of Units               1,800.00 A
Contribution amount          495,000.00 P=A*O
Total Fixed costs          247,500.00 K
Income before taxes          247,500.00 Q=P-K
Old Income before taxes          157,500.00 R
Increase in Income before taxes             90,000.00 S=Q-R
Increase by 57% T=S/R
The number is greater than 10% because increase sell price is the income to the company. There is no expense against it as the variable price is same. So the entire increase of 10% is a profit for the company.
Ans to 7
No. of Units to be sold               2,100.00 U
Sales       1,050,000.00 C*U
Variable manufacturing costs          367,500.00 E*U
Variable selling costs          210,000.00 G*U
Contribution          472,500.00 I*U
Total Fixed costs          247,500.00 K
Income before taxes          225,000.00
Taxes @ 40%             90,000.00
Income after taxes          135,000.00
Ans to 8
Income after taxes needed          120,000.00
Income before after taxes will be          200,000.00 120000/60% If tax rate is 40% then income after tax is 60% of income before tax through general mathematic calculation.
Add: Total fixed costs          247,500.00
Contribution needed          447,500.00 V
Contribution margin per ton                  225.00 I
Desired sales units               1,988.89 W=V/I
Desired sales value          994,444.44 X=W*C